Can you please help me with these simultaneous equations. I don't understand what I have to do. Thank you.
2x + y = 21
x - y = 6
and
u + v = 8
4u - v = 17

Answer by checkley75(3666)

X-Y=6 NOW FIND THE VALUE OF X OT Y THUS
X=Y+6 NOW SUBSTITUTE (Y+6) FOR X IN THE OTHERE EQUATION THUS:
2(Y+6)+Y=21 NOW SOLVE FOR Y
2Y+12+Y=21
3Y=21-12
3Y=9
Y=9/3
Y=3 ANSWER FOR X NOW REPLACE X BY 3 AND SOLVE FOR Y IN EITHER EQUATION
2X+3=21
2X=21-3
2X=18
X=18/2
X=9 ANSWER
PROOF
9-3=6
6=6
---------------------------------
U+V=8
U=8-V NOW SUBSTITUTE (8-V ) FOR U
4(8-V)-V=17
32-4V-V=17
-4V-V=17-32
-5V=15
V=15/-5
V=-3 ANSWER NOW SUBSTITUTE -3 FOR V & SOLVE FOR U
U+3=8
U=8-3
U=5 ANSWER
PROOF
4*5+(-3)=17
20-3=17
17=17